OpenStudy (lgbasallote):

obtain a family of solution for \[3(3x^2 + y^2)dx -2xydy = 0\] i just need this as basis for how the answer will look like but i still need steps

OpenStudy (cwtan):

sub y=vx and dy/dx=v+x dv/dx?

OpenStudy (cwtan):

I've 4gt how to integrate..... lol

OpenStudy (lgbasallote):

i think exact DE can be used though

OpenStudy (lgbasallote):

ugh it cant =_= isnt there any other way for this >.<

OpenStudy (cwtan):

This is what i have done: \[3(3x^2+y^2)dx=2xy dy\] \[\frac{dy}{dx}=\frac{3(3x^2+y^2)}{2xy}\] y=vx \[v+x\frac{dv}{dx}=\frac{3(3x^2+v^2x^2)}{2vx^2}\] \[v+x\frac{dv}{dx}=\frac{3(3+v^2)}{2v}\] \[x\frac{dv}{dx}=\frac{3(3+v^2)-2v^2}{2v}\] \[x\frac{dv}{dx}=\frac{9-v^2}{2v}\] \[\frac{1}{x}dx=\frac{2v}{9-v^2}dv\] ok following is integration which is the part i duno.........

OpenStudy (lgbasallote):

\[\ln x = -\ln |9 - v^2| + c\] \[\ln |x(9 - v^2)| = C\]

OpenStudy (binary3i):

x(9-v^2)=c x9-y^2/x =c

OpenStudy (lgbasallote):

\[\ln |x(\frac{9x^2 - y^2}{x^2})| = C\] \[\ln |\frac{9x^2 - y^2}{x}| = C\]

OpenStudy (lgbasallote):

\[\frac{9x^2 - y^2}{x} = C_1\] seems good to me..

OpenStudy (cwtan):

yea u've done it